Perdon Corporation manufactures safes—large mobile safes, and large walk-in stationary bank safes. As part of its annual budgeting process, Perdon is analyzing the profitability of its two products. Part of this analysis involves estimating the amount of overhead to be allocated to each product line. The information shown below relates to overhead.

Mobile Safes

Walk-in Safes

Units planned for production 200 50
Material moves per product line 300 200
Purchase orders per product line 450 350
Direct labor hours per product line 800 1,700

(a)

The total estimated manufacturing overhead was $260,000. Under traditional costing (which assigns overhead on the basis of direct labor hours), what amount of manufacturing overhead costs are assigned to: (Round answers to 2 decimal places, e.g. 12.25.)
(1) One mobile safe $ per unit
(2) One walk-in safe $ per unit
Answer #1

Manufacturing oh per labour unit

260000/(1700+800)=104

1. Manufacturing oh for one mobile safe

104*(800/2500)=33.28

2. Manufacturing oh for one walk in safe

104*(1700/2500)=70.72